Josie leaves home on her bicycle and rides out to the edge of the forest for three hours. on her return trip, along the same route, it takes her three-quarters of an hour longer. if her rate on the trip was two miles per hour slower than on the trip out forest, then
A) find the rate to bicycle INTO the forest?
B) what is the one-way distance from home to the forest?

Answer by ankor@dixie-net.com(22740)

Josie leaves home on her bicycle and rides out to the edge of the forest for three hours.
on her return trip, along the same route, it takes her three-quarters of an hour longer.
if her rate on the trip was two miles per hour slower than on the trip out forest, then
:
A) find the rate to bicycle INTO the forest?
let s = speed into the forest
then
(s-2) = speed returning
Write a distance equation, dist = time * speed
:
dist back = dist in
3.75((s-2) = 3s
3.75s - 7.5 = 3s
3.75s - 3 = 7.5
.75s = 7.5
s = 7.5/.75
s = 10 mph into the forest
:
B) what is the one-way distance from home to the forest?
dist = speed * time
d = 10 * 3
d = 30 miles
:
:
:
Check solutions using return trip
8 * 3.75 = 30